OpenStudy (anonymous):

3(y+7)/7y - (4y - 7)/6 = (7y+10)/3y - 2(7y-1)/21

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

(3y+21/7y)-(4y-7/6)=(7y+10/3y)-(14y-2/21) => (3y+21/7y)-(7y+10/3y)=-((14y-2/21)+(4y-7/6)) am working on the rest of the computation

OpenStudy (anonymous):

we will get i think -1596y²-567y+294=0 you need to do delta to find the solutions am not sure about the computation though

OpenStudy (anonymous):

ok so you should be able to simpliify it down to 42y^2 - 42y - 35 = 0 so factoring out a 7 would give you 6y^2 - 6y + 5 = 0 if you have learned imaginary roots: 0.5 +- i*(sqrt(21)/6)

OpenStudy (anonymous):

Move the right side of the equation over to the left side leaving zero on the right side and after simplifing the left side you will have the following equation: \[\frac{1}{42} \left(45-14 y-80 y^2\right)==0 \] Solve \[\left(45-14 y-80 y^2\right)==0 \]

OpenStudy (anonymous):

I haven't learned about imaginary roots yet. :/ I'm just learning algebraic fractions...so I'm realy confused

OpenStudy (anonymous):

\[y=\frac{1}{80} \left(-7\pm\sqrt{3649}\right)\]

OpenStudy (anonymous):

now im rly lost....

OpenStudy (anonymous):

After moving the right side to the left side, the left side should look like the following: \[\frac{1}{6}(7-4y)+\frac{3}{7}y(7+y)-\left(-\frac{2}{21}(-1+7y)+\frac{1}{3}y(10+7y)\right)=0\] Get rid of the fractions by multipling each fraction in front of it's term by 42, the smallest number divisible by all of the fraction denominators. Do the multiplications and collect like powers of x.